solve radical equation and check all proposed solutions
square root ofx+10 = x+4

Found 2 solutions by amarjeeth123, stanbon:
Answer by amarjeeth123(569)

sqrt(x+10)=x+4
Squaring on both sides
(x+10)=(x+4)^2
x+10=x^2+8x+16
x^2+7x+6=0
x^2+6x+x+6=0
x(x+6)+1(x+6)=0
(x+1)(x+6)=0
x+1=0 or x+6=0
x=-1 or x=-6

Answer by stanbon(75887)

square root ofx+10 = x+4
sqrt(x+10) = x+4
------
Square both side sides to get::
x+10 = x^2 + 8x + 16
--------
x^2 + 7x + 6 = 0
-------
Factor::
(x+6)(x+1) = 0
-----
x = -1 or x = -6
-----
Checking::
sqrt(x+10) = x+4
Check x = -1::
sqrt(9) = -1+4
3 = 3
=================
Check x = -6
sqrt(4) = -2
False
----------------
Ans:: x = -1
--------------------------
Cheers,
Stan H.